OpenStudy (anonymous):

Find the missing numbers. cost =350 Selling price N/A Markup on =cost Markup Rate=70% Markup Amount N/A MY answer : Selling Price = $947.00, Markup Amount = $478.00

OpenStudy (radar):

Your mark-up seems hight, are you sure you used the correct rate of 70%?

OpenStudy (anonymous):

ok can you show me what i did wrong

OpenStudy (radar):

Cost = $350 Mark-up rate is 70% mark-up= .70 X 350 = ?????

OpenStudy (radar):

Did you multiply cost by mark-up rate???

OpenStudy (anonymous):

thank you

OpenStudy (radar):

what did you get for the markup? I get $245

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got 245

OpenStudy (anonymous):

do you think you can help me with my assignment if i post it

OpenStudy (radar):

as long as it is not a long convoluted assignment, I ma be of help.

OpenStudy (radar):

Which assignment? The triples or the other one?

OpenStudy (anonymous):

both if possible

OpenStudy (radar):

The triples has been provided to you earlier. I don't think I could add anything more.

OpenStudy (anonymous):

i did not understand it

OpenStudy (anonymous):

so if you could help me understand it i would be very thankful

OpenStudy (radar):

Let me look at wiki and see if I can figure out something (like the assignment suggested)

OpenStudy (anonymous):

k

OpenStudy (radar):

I looked at the wikipedia reference provided in your assignment, and it provided several triple generators. The one that seems the simplest is this one: a=2n+1 b=2n(n+1) c=2n(n+1) +1 Don't you agree?

OpenStudy (radar):

You can use this to create triples. like letting n=1, you will get a=3, b=4, and c=5 And according to the pythagorean theorem \[3^{2}+4^{2}=5^{2}\]

OpenStudy (anonymous):

but how

OpenStudy (radar):

if you use n=2 you will get: a=5 b=12 c=13 \[5^{2}+12^{2}=13^{2}\]\[25+144 =169 \]yup you can proceed with your home workl.

OpenStudy (anonymous):

hang on

OpenStudy (anonymous):

so my answer would be put in how

OpenStudy (anonymous):

that would be me showing my work

OpenStudy (radar):

Follow the assignment template. You are to generate a total of 5 triples, using a triple generator. I have provided you a triple generator. I have already done two triples and used the pythagorean theorem to prove the triples, you must do the other three. Don't you see how this is done?

OpenStudy (anonymous):

for the first one that you have done for me

OpenStudy (anonymous):

would this be the second part

OpenStudy (anonymous):

9+16=25

OpenStudy (radar):

Don't forget to write an intro paragraph. in that you could mention the source of the triple wikipedia. And that was the first one not the second (my mistake)

OpenStudy (anonymous):

but what im asking is [3^2 +4^2=5^2 \] comes out to 9+16=25

OpenStudy (radar):

Possibily in the intro you could state you are going to use the right triangle as the pythagoras example to test and demonstrate your triples.

OpenStudy (radar):

Yes 3 squared is 9, 4 squared is 16, 9+16 =25 which is 5 squared. This is the test for your triple. When you generate your triple show the work like: (for n=1) the first one "a=2(1)+1=3" "b=2(1)(1+1)=4" "c=2(1)(1+1)+1=5"

OpenStudy (radar):

Do the remaining 4, then do the summarizing paragraph.

OpenStudy (radar):

we have used n=1 and n=2 for thefirst two, use n=3, 4, 5 for the remaining 3.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

what would a b and c be for all three still not clicking

OpenStudy (radar):

Good luck with this. I don't know what kind of course you are taking, but it does have some interesting problems with "interesting" protocols for the answers lol. Bye